Make New Entry,
Make Followup Entry
User name Sirish
Log entry time 15:47:53 on November12,2008
Entry number 248409
Compton chicane beam tune has been successfuly completed.
Difficulties were encountered in threading the beam through the newly installed e-detector chamber and beam pipes. The beam was scraping near the e-detector and causing ep-ion chamber trips.
The orbit of the beam near the e-detector had to be raised by 1.5 mm from before for clear passage of the beam.
Now that a new beam tune through the Compton chicane has been established, stability of the tune should be monitored over the next day or so.